OpenStudy (inowalst):

Can someone explain how to change radical equivalents to rational exponents and vice versa?

OpenStudy (unklerhaukus):

\[\large x^{1/q}=\sqrt[q]x\\ \large x^{p/q}=\sqrt[q]x^p\] For example \[64^{1/3}=\sqrt[3]{64}=\sqrt[3]{4^3}=4\] \[4^{3/2}=\sqrt[2]{4^3}=\sqrt{4^3}=\sqrt{64}=\sqrt{8^2}=8\]

OpenStudy (inowalst):

Okay. What change was that? Radical to rational or the other way around?

OpenStudy (anonymous):

Radicals are \[\sqrt{x}\]Rationals are \[x ^{1/3}\]

OpenStudy (inowalst):

Ohhh okay. Thank you! :D

OpenStudy (theraggedydoctor):

dot.